# A Systematic Survey on Large Language Models for Algorithm Design

### v2

@@ 6.4. Industry
To enhance the reliability and availability of cloud services, conducting root cause analysis (RCA) for cloud incidents is essential. RCACopilot (Chen et al., 2024b) effectively matches incoming incidents with the appropriate incident handlers based on their alert types. It aggregates critical runtime diagnostic information, predicts the incident’s root cause category, and provides explanatory results without the need for manual investigations.
More applicable scenarios include various industrial design challenges, such as chip design (Chang et al., 2024), car shape design (Wong et al., 2024b), aircraft concept design (Lovaco et al., 2024) UI design (Feng et al., 2023a), and robot design (Yu et al., 2023). LLMs primarily leverage billions of web resources. However, practical applications require fine-tuning with thousands of specific representations of design intent to recognize the distinct patterns associated with each design task.
Other traditional industrial tasks, such as itinerary planning, can also be enhanced by leveraging the power of LLMs (Xie et al., 2024). LLM-based open-domain urban itinerary planning (Tang et al., 2024b) combines spatial optimization with LLMs to create detailed urban itineraries tailored to users’ requests. This approach extends the scope of itinerary planning by generating personalized urban itineraries based on user requests in natural language.
@@ 7.1. Scalability
One of the primary limitations of LLMs in algorithm design is their scalability, particularly for handling long inputs and outputs. LLMs are limited by a fixed context size, restricting the amount of information they can process at once, which is a challenge for complex algorithmic tasks requiring detailed specifications and extensive content. Additionally, even when input and output lengths are sufficient, LLMs struggle with comprehension and reasoning over long inputs, impacting their effectiveness in designing complex algorithms. For instance, LLMs tend to solve only low-dimensional problems when used as optimizers (Yang et al., 2024), and they are employed to design a function or heuristic component of an algorithm rather than the entire algorithm when used as designers (Liu et al., 2024d).
@@ 7.3. Security
Integrating LLMs into algorithm design introduces significant security risks, including the mishandling of personal data and the potential creation of harmful code. LLMs are trained on extensive datasets that often contain sensitive information, posing a risk of privacy breaches if these models inadvertently generate outputs influenced by this data. Additionally, the autonomous nature of LLMs can lead to the unintentional development of flawed or malicious algorithms, a critical concern in sectors like financial services and personal data management (Zhao et al., 2024). To address these issues, it is vital to establish stringent data governance and model training protocols that prevent the inclusion of sensitive data, rigorously test algorithms under various conditions to prevent unintended behaviors, and continuously update security measures to counteract emerging vulnerabilities.
@@ 7.4. Cost
The utilization of LLMs in algorithm design also faces challenges related to costs, both in terms of model training and usage. Training domain LLMs requires significant computational resources, often involving extensive GPU or TPU usage over several weeks or even months, which incurs high financial and time costs. Once trained, the inference cost—the computational expense involved in applying the model to algorithm design—also becomes a critical factor, especially when the model is used frequently or requires rapid responses, such as the evolutionary search used in many LLM4AD works (Romera-Paredes et al., 2024). For practical applications in algorithm design, where multiple iterations and tests might be necessary, these costs can accumulate, making the use of LLMs potentially less viable for researchers without substantial computational budgets.
@@ 7.5. Idea Innovation
Despite the numerous works on LLM4AD, their ability to produce truly novel ideas is still under scrutiny. LLMs excel at interpolation within the range of their training data but may struggle with extrapolation to new, unseen scenarios. In the context of algorithm design, this means that while LLMs can effectively suggest algorithms or solutions that closely resemble or incrementally improve upon existing ones, they may not be as effective at pioneering radically new algorithmic concepts or methods. The reliance on large volumes of training data also raises concerns about the model’s dependency on historical data, which may bias its outputs towards conventional methods and potentially stifle innovation. While LLM does show the capability of generating new ideas in some tasks (Yao et al., 2024a), how to continuously explore practical novel ideas remains a challenge (Hu et al., 2024).
@@ 8.4. LLM-based Algorithm Assessment
LLMs can be helpful in algorithm assessment. Some attempts have been carried out to automatically evaluate algorithms and assess algorithm design. For example, Sterbini and Temperini (2024) show that LLMs can be used to analyze algorithm descriptions and compare them to the task description given. Such text similarity correlates well both with the grades given by the teacher and with the grades given by the students. OMNI-EPIC (Faldor et al., 2024) utilizes foundation models to automatically produce code that defines the next engaging and learnable tasks. By generating environments and reward functions, OMNI-EPIC has the capability to create a wide range of simulated learning tasks for algorithm evaluation. We expect more studies on LLM-based algorithm assessment.
@@ 8.6. Fully Automatic Algorithm Design
Fully automatic algorithm design faces two primary challenges: 1) the generation of novel algorithmic ideas and 2) the creation of complex, lengthy code. While the generation of new ideas has been investigated in some works (Liu et al., 2024d), the complete design of algorithms (instead of a heuristic component), including both heuristic components and detailed code implementations, remains a challenge. Existing applications typically focus on automating components within predefined algorithm frameworks rather than creating new algorithms from scratch. Future research needs to tackle these complexities to advance the field of fully automatic algorithm design.
@@ 9. Conclusion
This paper has provided a systematic and up-to-date survey on large language models for algorithm design (LLM4AD). By systematically reviewing a significant corpus of main contributions in this emerging research field, this paper not only highlights the current state and evolution of LLM applications in algorithm design but also introduces a novel multi-dimensional taxonomy with comprehensive lists of papers that categorize the roles and functionalities of LLMs in this domain. This taxonomy serves as a framework for both academic and industrial researchers to understand and navigate the complex landscape of algorithm design using LLMs. We have also discussed the limitations and challenges currently faced by the field, such as issues of scalability, interpretability, and security. Moreover, the proposed future research directions aim to inspire and guide subsequent studies.
As we look forward, it is clear that the intersection of LLMs and algorithm design holds promising potential for revolutionizing how algorithms are developed and implemented. The ability of LLMs to automate and innovate within the algorithm design process could lead to more efficient, effective, and creative algorithms for complex problems across various domains. We hope this paper can contribute to a deeper understanding of this potential and set the stage for future innovations and collaborations in this promising avenue of research.

### v3

@@ 6.4. Industry
To enhance the reliability and availability of cloud services, conducting root cause analysis (RCA) for cloud incidents is essential. RCACopilot (Chen et al., 2024c) effectively matches incoming incidents with the appropriate incident handlers based on their alert types. It aggregates critical runtime diagnostic information, predicts the incident’s root cause category, and provides explanatory results without the need for manual investigations.
More applicable scenarios include various industrial design challenges, such as chip design (Chang et al., 2024), car shape design (Wong et al., 2024b), aircraft concept design (Lovaco et al., 2024) UI design (Feng et al., 2023a), and robot design (Yu et al., 2023). LLMs primarily leverage billions of web resources. However, practical applications require fine-tuning with thousands of specific representations of design intent to recognize the distinct patterns associated with each design task. Other traditional industrial tasks, such as itinerary planning, can also be enhanced by leveraging the power of LLMs (Xie et al., 2024). LLM-based open-domain urban itinerary planning (Tang et al., 2024b) combines spatial optimization with LLMs to create detailed urban itineraries tailored to users’ requests.
@@ 7.1. Scalability
One of the primary limitations of LLMs in algorithm design is their scalability. LLMs are limited by a fixed context size, restricting the amount of information they can process at once, which is a challenge for complex algorithmic tasks requiring detailed specifications and extensive content. Additionally, even when input and output lengths are sufficient, LLMs struggle with comprehension and reasoning over long inputs, impacting their effectiveness in designing complex algorithms. For instance, LLMs tend to solve only low-dimensional problems when used as optimizers (Yang et al., 2024), and they are employed to design a function or heuristic component of an algorithm rather than the entire algorithm when used as designers (Liu et al., 2024e).
@@ 7.3. Security
Integrating LLMs into algorithm design introduces significant security risks, including the mishandling of personal data and the potential creation of harmful code. LLMs are trained on extensive datasets that often contain sensitive information, posing a risk of privacy breaches if these models inadvertently generate outputs influenced by this data. Additionally, the autonomous nature of LLMs can lead to the unintentional development of flawed or malicious algorithms, which is a critical concern in sectors like financial services and personal data management (Zhao et al., 2024). To address these issues, it is vital to establish stringent data governance and model training protocols that prevent the inclusion of sensitive data, rigorously test algorithms under various conditions to prevent unintended behaviors, and continuously update security measures to counteract emerging vulnerabilities.
@@ 7.4. Cost
The utilization of LLMs in algorithm design also faces challenges related to high cost in model training and usage. Training domain LLMs requires significant computational resources, often involving extensive GPU or TPU usage over several weeks or even months, which incurs high financial and time costs. Once trained, the inference cost—the computational expense involved in applying the model to algorithm design—also becomes a critical factor, especially when the model is used frequently, such as the evolutionary search used in many LLM4AD works (Romera-Paredes et al., 2024). For practical applications in algorithm design, where multiple iterations and tests might be necessary, these costs can accumulate, making the use of LLMs potentially less viable for researchers without substantial computational budgets.
@@ 7.5. Idea Innovation
Despite extensive research on LLM4AD, their capacity for generating novel ideas remains questionable. LLMs are adept at interpolating within their training data but often falter in extrapolating to new scenarios. In algorithm design, this means LLMs can suggest or slightly enhance existing algorithms but struggle to develop radically new methods. Although there is evidence of LLMs producing new ideas in certain tasks (Yao et al., 2024a), consistently generating practical, innovative concepts continues to be a challenge (Hu et al., 2024).
@@ 8.3. Interaction with Human Experts
8.4. Multi-objective Algorithm Design
There are usually multiple criteria (e.g., optimal solution generated by algorithm, optimization efficiency, and generalization performance) in real-world algorithm design. The majority of existing LLM4AD papers consider one objective (Liu et al., 2024e; Romera-Paredes et al., 2024) focusing only the optimal solutions. A preliminary attempt has been conducted by Yao et al. (2024b), where a dominance-dissimilarity measurement, considering both the dominance relation in objective space and the distance in code space, is designed for efficient multi-objective search in the heuristics space. More future works are anticipated to develop new methods for effective multi-objective algorithm design.
@@ 8.5. LLM-based Algorithm Assessment
LLMs can be benefitial in algorithm assessment. Some attempts have been carried out on test instance generation and algorithm evaluation. For example, Jorgensen et al. (2024) use LLMs to generate test cases of increasing difficulties to enhance genetic programming and OMNI-EPIC (Faldor et al., 2024) utilizes LLMs to automatically produce code that defines the next engaging and learnable tasks for algorithm evaluation. Further research in this area is anticipated.
@@ 8.7. Fully Automatic Algorithm Design
Fully automatic algorithm design faces two primary challenges: 1) the generation of novel algorithmic ideas and 2) the creation of complex, lengthy code. While the generation of new ideas has been investigated in some works (Liu et al., 2024e), the complete design of algorithms (instead of a heuristic component), remains a challenge. Existing applications typically focus on automating components within predefined algorithm frameworks rather than creating new algorithms from scratch. Future research needs to tackle these complexities to advance the field of fully automatic algorithm design.
@@ 9. Conclusion
This paper has provided a systematic and up-to-date survey on large language models for algorithm design (LLM4AD). By systematically reviewing a significant corpus of main contributions in this emerging research field, this paper not only highlights the current state and evolution of LLM applications in algorithm design but also introduces a multi-dimensional taxonomy with comprehensive lists of papers that categorize the roles and functionalities of LLMs in this domain. This taxonomy serves as a framework for both academic and industrial researchers to understand and navigate the complex landscape of algorithm design using LLMs. We have also identified the limitations and challenges currently faced by the field, such as issues of scalability, interpretability, and security. Moreover, we have highlighted and discussed some promising research directions to inspire and guide future studies.
As we look forward, the intersection of LLMs and algorithm design holds promising potential for revolutionizing how algorithms are developed and implemented. We hope this paper can contribute to a deeper understanding of this potential and set the stage for future innovations and collaborations in this emerging avenue of research.